Determine the number of solutions to the given system without graphing.
y=4x+5
20x-5y=4

Answer by Fombitz(32388)

From eq. 1,
4x-y=-5
20x-5y=-25
Compare to eq. 2,
20x-5y=4
Which leads to
-25=4
So there is no solution.
The two lines are parallel.